OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 + 24x + 62? Answer x = 6 x = –6 x = –10 x = 10

OpenStudy (anonymous):

= 2(x^2 + 12x + 31) = 2(x+ 6)^2 = 31 - 36 x = -6